OpenStudy (anonymous):

what is the solution for this (a) The length of a fence required to enclose a rectangular fence is 3000 metres. (i) Write a system of equations to find the dimensions of the field if, the difference between its length and width is 50 metres. (ii) Find the dimensions of the field.

OpenStudy (anonymous):

Let length be x then width will be x-50 . The perimeter of the rectangle is given by 2(l+b) = 3000 , now substitute the values for l= x and b= x-50 and find the value of x

OpenStudy (anonymous):

thanku so much nishika1 7

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

\[\large \color{green}{Perimeter_{rectangle} = 2 \times Length + 2 \times Width = 2(L + W)}\]

OpenStudy (anonymous):

thanks waterineyes

OpenStudy (anonymous):

so the ans will be l=1550

OpenStudy (anonymous):

nishika1 7 correct me if I'm wrong